OpenStudy (anonymous):

What are the allowable forms of an expression for trigonometric substitution? I know that \sqrt{x^2 - a^2} \sqrt{a^2 - x^2} \sqrt{x^2 + a^2} ^- these count, but what are others?

OpenStudy (anonymous):

ArcSin=> 1/root(1-u^2) (du/dx) ArcTan=> 1/(1+u^2) (du/dx) ArcCos=> -1/root(1-u^2) (du/dx) ArcCot=> -1/(1+u^2) (du/dx)

OpenStudy (turingtest):

\[\sqrt{a^2-b^2x^2}\implies x=\frac ab\sin\theta\]\[\sqrt{b^2x^2-a^2}\implies x=\frac ab\sec\theta\]\[\sqrt{b^2x^2+a^2}\implies x=\frac ab\tan\theta\]those are the only ones you really need to know

OpenStudy (anonymous):

Does a^2 + x^2 <-- this count?

OpenStudy (turingtest):

yes, that is my third formula above with b=1